OpenStudy (anonymous):

A girl dilereing newspapers covers her route by traveling 3.00 blocks west, 4.00 blocks north, and then 6.00 blocks east. A) What is her resultant displacement? B) What is the total distance she travels?

OpenStudy (anonymous):

b

OpenStudy (johnweldon1993):

Well she went 3 blocks west....4 north...and then 6 east...so displacement would be  Distance between those 2 point = \(\large \sqrt{(3)^2 + (4)^2} = \sqrt{25} = 5\) so she is 5 blocks from where she started The TOTAL distance she traveled would be 3 + 4 + 6 = 13 blocks

OpenStudy (anonymous):

Ok but what is the resultant displacement

OpenStudy (anonymous):

ok let me thin first

OpenStudy (johnweldon1993):

It would be the 5 I posted before the 13 :) The displacement would be how far from start you are...well she ends up at that 2nd point...so all you need to do is find the distance from start to finish...which can be done using pythagorea's theorem...which came out to 5

OpenStudy (anonymous):

ok thanks